OpenStudy (anonymous):

WILL FAN AND GIVE MEDAL ALGEBRA 2 What are the zeros of the function and their multiplicity? f(x)=x^4+x^3-6x^2 Show all work.

Directrix (directrix):

Solve this: x^4 +x^3 - 6x^2 = 0 Begin by factoring out x^2 By that, I mean this: x^4 +x^3 - 6x^2 = x^2 ( ? + ? + ?) = 0 @WesternBurger

OpenStudy (anonymous):

x^2(x^2+x+6) right?

OpenStudy (anonymous):

x^2(x^2+x-6) I meant. And then... x^2(x+3)(x-2) so.. Are the zeros -3, 2? And multiplicity 1 for both of them?

OpenStudy (anonymous):

@Directrix can you check my work please? I have a lot more questions to get through.

OpenStudy (anonymous):

I just don't get what to do with the x^2 part.

Directrix (directrix):

x^2(x+3)(x-2) = 0 Use the Zero Product Property x^2 = 0 OR x + 3 = 0 OR x-2 = 0

Directrix (directrix):

Are the zeros -3, 2? And multiplicity 1 for both of them? YES, Those are two of the four roots. The multiplicities are correct.

Directrix (directrix):

But, what about the x^2 = 0. From it comes x = 0 or x = 0 which is a double root.

OpenStudy (anonymous):

Ohhhhh ok. So the answer is: Zeros: x=-3(mult. 1), 2(mult. 1) and 0(mult. 2) Correct?

Directrix (directrix):

Correct.

OpenStudy (anonymous):

Thank you so much! Just one quick thing... I have a few more questions I was wondering if you could help me with? If you have the time.

Directrix (directrix):

The graph is tangent to the x axis at 0,0

OpenStudy (anonymous):

Oh I'm not on the whole graph stuff yet lol.

Directrix (directrix):

You can close this thread and post them one per thread. A gazillion people are asking me to help but you do respond to questions so I'll try to help you.

OpenStudy (anonymous):

Ok, so I can just tag you or something?

Directrix (directrix):

Yes.

OpenStudy (anonymous):

Ok, thank you :)